OpenStudy (anonymous):

[x^2-9]/[x-2] HA:=3 VA:= -3 OA:= 2 Are these correct?

OpenStudy (lgbasallote):

what's OA? original asymptote?

OpenStudy (anonymous):

oblique

OpenStudy (lgbasallote):

hmm one slight mistake...VA should be 2

OpenStudy (lgbasallote):

because you equate x - 2 to 0 then solve for x so x = 2

OpenStudy (lgbasallote):

assuming VA is vertical asymptote

OpenStudy (anonymous):

ty. ahh very good x^2-16/6x-24 disct = x can not = 4 is that correct?

OpenStudy (lgbasallote):

disct means discriminant?

OpenStudy (anonymous):

discontinuities

OpenStudy (lgbasallote):

oh lol

OpenStudy (lgbasallote):

yes x = 4 is a removable discontinuity

OpenStudy (anonymous):

so it can not = 4?

OpenStudy (anonymous):

or it does

OpenStudy (anonymous):

but can not = -4

OpenStudy (anonymous):

[x^2-x-12]/[x-4] disc.= -3 x int=(3,0) y int= (0,-3) shape= straight line are these correct?